# Transition Constrained Bayesian Optimization via Markov Decision Processes

### v2

@@ 1 Introduction
Many areas in the natural sciences and engineering deal with optimizing expensive black-box functions. Bayesian optimization (BayesOpt) [16, 29, 52], a method to optimize these problems using a probabilistic surrogate, has been successfully applied to myriad examples, e.g. tuning the hyper-parameters of expensive-to-train models [24], robotics [34], battery design [14], and drug discovery [6]. However, state-of-the-art algorithms are often ill-suited when physical sciences interact with potentially dynamic systems. In such circumstances, real-life constraints limit our future decisions while depending on the prior state of our interaction with the system. This work focuses on transition constraints influencing future choices depending on the current state of the experiment. In other words, reaching certain parts of the decision space (search space) requires long-term planning in our optimization campaign. This effectively means we address a general sequential-decision problem akin to those studied in reinforcement learning (RL) or optimal control for the task of optimization. We assume the transition constraints are known a priori to the optimizer.
Applications with transition constraints include chemical reaction optimization [57, 13], environmental monitoring [4, 39], lake surveillance with drones [23, 19, 49], energy systems [45], vapor compression systems [41], electron-laser tuning [31] and seabed identification [55]. For example, Figure 1 depicts an application in environmental monitoring where autonomous sensing vehicles must avoid obstacles (similar to Hitz et al. [23]). Our main focus application are transient flow reactors [35, 50]. Such reactors allow efficient data collection by obtaining semi-continuous time-series data rather than a single measurement after reaching the steady state of the reactor. As we can only change the inputs of the reactor continuously and slowly to maintain quasi-steady-state operation, allowing arbitrary changes, as in conventional BayesOpt, would result in measurement sequences which are not possible due to physical limitations.
Problem Statement. More formally, we design an algorithm to identify the optimal configuration of a physical system governed by a black box function $f$, namely, $x^{\star}=\arg\max_{x\in\mathcal{X}}f(x).$ The set $\mathcal{X}$ summarizes all possible system configurations, the so called search space. We assume that we can sequentially evaluate the unknown function at specific points $x$ in the search space and obtain noisy observations, $y=f(x)+\epsilon(x)$, where $\epsilon$ has a known Gaussian likelihood, which is possibly heteroscedastic. We assume that $f$ can be modelled probabilistically using a Gaussian process prior that we introduce later. Importantly, the order of the evaluations is dictated by known, potentially stochastic, dynamics modelled by a Markov chain that limits our choices of $x\in\mathcal{X}$.
@@
Figure 1: Representative task: finding pollution in a river while following the current. (a) Problem formulation: The star represents the maximizer and the arrows the Markov dynamics. (b) Objective formulation: Orange balls represent potential maximizers, with size corresponding to model uncertainty. (c) Optimization: Deploy a potentially stochastic policy that minimizes our objective.
@@ 2 Background
We first review relevant background from experimental design, BayesOpt, and RL that we build upon.
@@ 2.1 Maximum Identification: Experiment Design Goal
Suppose we are in step $t$ out of $T$, then let $\mathbf{X}_{t}$ be the set of previous queries, we seek identify new $\mathbf{X}_{\text{new}}$ that when evaluated minimize the probability of returning a sub-optimal arm at the end. For a given function draw $f$, the probability of returning a wrong maximizer $z\neq x_{f}^{\star}$ is $P(\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\geq 0|f)$. We can then consider the worst-case probability across potential maximizers, and taking expectation over $f$ we obtain a utility through a tight upper-bound on it:
$\min_{\mathbf{X}_{\text{new}}}\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}P(\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\geq 0|f)\right]\stackrel{{\scriptstyle\cdot}}{{\leq}}\min_{\mathbf{X}_{\text{new}}}\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\frac{k_{\mathbf{X}_{t}\cup\mathbf{X}_{new}}(z,x_{f}^{\star})}{(f(z)-f(x_{f}^{\star}))^{2}}\right]$ (3)
The expectation is on the current prior (posterior up to $\mathbf{X}_{t}$), the kernel $k$ is the posterior kernel given observations $\mathbf{X}_{t}\cup\mathbf{X}_{\text{new}}$. Since we consider the probability of an error, it is more appropriate to talk about minimizing instead of ‘maximizing the utility’ but the treatment is analogous. The non-trivial distribution of $f(x^{\star})$ [21] renders the utility intractable; therefore we employ a simple and tractable upper bound on the objective (3) by minimizing the uncertainty among all pairs in $\mathcal{Z}$:
@@
Utility with Embeddings. For illustrative purposes, consider a special case where the kernel $k$ has a low rank due to existence of embeddings $\Phi(x)\in\mathbb{R}^{m}$, i.e., $k(x,y)=\Phi(x)^{\top}\Phi(y)$. Such embeddings can be, e.g., Nyström features [59] or Fourier features [36, 44]. While not necessary, these formulations make the objectives considered in this work more tractable and easier to expose to the reader. With the finite rank assumption, the random function $f$ becomes,
@@ 2.2 Markov Decision Processes
To model the transition constraints, we use the versatile model of Markov Decision processes (MDPs). We assume an environment with state space $\mathcal{X}$ and action space $\mathcal{A}$, where we interact with an unknown function $f:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$ by rolling out a policy for $H$ time-steps (horizon) and obtain a trajectory, $\tau=(x_{0},a_{0},x_{1},a_{1},...,x_{H-1},a_{H-1})$. From the trajectory, we obtain a sequence of noisy observations $y(\tau):=\{y(x_{0},a_{0}),...,y(x_{H-1},a_{H-1})\}$ s.t. $y(x_{h})=f(x_{h},a_{h})+\epsilon(x_{h},a_{h})$, where $\epsilon(x_{h},a_{h})$ is zero-mean Gaussian with known variance which is potentially state and action dependent. The trajectory is generated using a known transition operator $P(x_{h+1}|x_{h},a_{h})$. A Markov policy $\pi(a_{h}|x_{h})$ is a mapping that dictates the probability of action $a_{h}$ in state $h$. Hence, the state-to-state transitions are $P(x_{h+1},x_{h})=\sum_{a\in\mathcal{A}}\pi_{h}(a|x_{h})P(x_{h+1}|x_{h},a)$. In fact, an equivalent description of any Markov policy $\pi$ is its visitation of states and actions, which we denote $d\in\mathcal{D}$, where
@@ 2.3 Experiment Design in Markov Chains
Notice that the utility $U$ in Eq. 6 depends on the states visited and hence states of the trajectory. In our notation, $\mathbf{X}_{t}$ will now form a set of executed trajectories. With deterministic dynamics, we could optimize over trajectories, but this would lead to an exponential blowup (i.e. $|S|^{H}$). In fact, for stochastic transitions, we cannot pick the trajectories directly and the utility would be random $U(\mathbf{X}_{t})$. Hence, we focus on expected utility over the randomness of the policy and the environment, namely,
$\mathcal{U}(d_{\pi}):=U(\mathbb{E}_{\tau_{1}\sim\pi_{1},\dots\tau_{t}\sim\pi_{t}}[\mathbf{X}_{t}]).$ (7)
This formulation stems from Mutny et al. [39] who try to tractably solve such objectives that arise in experiment design by performing planning in MDPs. They focus on learning linear operators of an unknown function, unlike identifying a maximum, as we do here. The key observation they make is that any policy $\pi$ induces a distribution over the state-action visitations, $d_{\pi}$. Therefore we can reformulate the problem of finding the optimal policy, into finding the optimal distribution over state-action visitations as: $\min_{d_{\pi}\in\mathcal{D}}\mathcal{U}(d_{\pi})$, and then construct policy $\pi$ via marginalization. We refer to this optimization as the planning problem. The constraint $\mathcal{D}$ encodes the dynamics of the MDP.
@@ 2.4 Additional Related Works
The most relevant prior work to ours is exploration in reinforcement learning through the use of Markov decision processes as in Mutny et al. [39] and convex reinforcement learning of Hazan et al. [20], Zahavy et al. [64] which we will use to optimize the objective. Other related works are:
Optimizing over sequences. Previous work has focused on planning experimental sequences for minimizing switching costs [13, 45, 61] however they are only able adhere to strict constraints under truncation heuristics [49, 15, 41]. Concurrent work of Che et al. [8] tackles a constrained variant of a similar problem using model predictive control with a different goal.
@@
Non-myopic Bayesian Optimization Look-ahead BayesOpt [18, 32, 27, 33, 40, 9] seeks to improve the greedy aspect of BayesOpt. Such works also use an MDP problem formulation, however, they define the state space to include all past observations (e.g. [28, 1]). This comes at the cost of simulating expensive integrals, and the complexity grows exponentially with the number of look-ahead steps (usually less than three steps). Our work follows a different path, we maintain the greedy approach to control computational efficiency (i.e. by optimizing over the space of Markovian policies), and maintain provable and state-of-art performance. Even though the optimal policy through non-myopic analysis is non-Markovian, in Sec. 4, we show that adaptive resampling iteratively approximates this non-myoptic optimal policies in a numerically tractable way via receeding horizon planning. In our experiments we comfortably plan for over a hundred steps.
@@ 3 Transition Constrained BayesOpt
Motivated by our practical experiments with chemical reactors, we distinguish two different types of feedback. With episodic feedback we can be split the optimization into episodes. At the end of each episode of length $H$, we obtain the whole set of noisy observations. On the other hand, instant feedback is the setting where we obtain a noisy observation immediately after querying the function. Asynchronous feedback describes a mix of previous two, where we obtain observations with unspecific a delay.
@@ 3.1 Expected Utility for Maximizer Identification
where $\mathbf{V}(d_{\pi})=\left(\sum_{x,a\in\mathcal{X}\times\mathcal{A}}\frac{d_{\pi}(x,a)\Phi(x,a)\Phi(x,a)^{\top}}{\sigma^{2}(x,a)}+\mathbf{I}\right)$. The variable $d_{\pi}(x,a)$ is a state-action visitation, $\Phi(x)$ are e.g. Nyström features of the GP. We prove that the function is additive in terms of state-action pairs in Lemma D.1 in Appendix D, a condition required for the expression as a function of state-action visitations [39]. Additionally, by rewriting the objective in this form, the dependence and convexity with respect to the state-action density $d_{\pi}$ becomes clear as it is only composition of a linear function with an inverse operator. Also, notice that the constraint set is a convex polytope. Therefore we are able to use convex optimization to solve the planning problem (see Sec. 4).
@@
In the Eq. (8) we need to approximate a maximum over all input pairs in $\mathcal{Z}$. While this can be enumerated in the discrete case without issues, it poses a non-trivial constrained optimization problem when $\mathcal{X}$ is continuous. As an alternative, we propose approximating the set $\mathcal{Z}$ using a finite approximation of size $K$ which can be built using Thompson Sampling [30] or through maximization of different UCBs for higher exploitation (see Appendix E.1). In Appendix E.5, we numerically benchmark reasonable choices of $K$, and show that the performance is not significantly affected by them.
@@
  Input: Procedure for estimating sets of maximizers, initial point $x_{0}$, initial set of maximizers $\mathcal{Z}_{0}$
@@
           $y_{t,h+1}=f(x_{t,h+1})+\epsilon_{t,h};\ \mathcal{GP}_{t,h},\ \mathcal{Z}_{t,h}\leftarrow\text{Update}(\mathbf{X}_{t,h},\mathbf{Y}_{t,h})$ // update model
@@
        $\mathbf{Y}_{t,H}=f(\mathbf{X}_{t,H})+\vec{\epsilon}_{t,:};\ \mathcal{GP}_{t+1,:},\ \mathcal{Z}_{t+1,:}\leftarrow\text{Update}(\mathbf{X}_{t,H},\mathbf{Y}_{t,H})$ // update model

### v3

@@ 1 Introduction
Many areas in the natural sciences and engineering deal with optimizing expensive black-box functions. Bayesian optimization (BayesOpt) [1, 2, 3], a method to optimize these problems using a probabilistic surrogate, has been successfully applied to a myriad of examples, e.g. hyper-parameter selection [4], robotics [5], battery design [6], laboratory equipment tuning [7], and drug discovery [8]. However, state-of-the-art algorithms are often ill-suited when physical sciences interact with potentially dynamic systems [9]. In such circumstances, real-life constraints limit our future decisions while depending on the prior state of our interaction with the system. This work focuses on transition constraints influencing future choices depending on the current state of the experiment. In other words, reaching certain parts of the decision space (search space) requires long-term planning in our optimization campaign. This effectively means we address a general sequential-decision problem akin to those studied in reinforcement learning (RL) or optimal control for the task of optimization. We assume the transition constraints are known a priori to the optimizer.
Applications with transition constraints include chemical reaction optimization [10, 11, 12], environmental monitoring [13, 14, 15, 16, 17], lake surveillance with drones [18, 19, 20], energy systems [21], vapor compression systems [22], electron-laser tuning [23] and seabed identification [24]. For example, Figure 1 depicts an application in environmental monitoring where autonomous sensing vehicles must avoid obstacles (similar to Hitz et al. [18]). Our main focus application are transient flow reactors [25, 26, 27]. Such reactors allow efficient data collection by obtaining semi-continuous time-series data rather than a single measurement after reaching the steady state of the reactor. As we can only change the inputs of the reactor continuously and slowly to maintain quasi-steady-state operation, allowing arbitrary changes, as in conventional BayesOpt, would result in measurement sequences which are not possible due to physical limitations.
Problem Statement. More formally, we design an algorithm to identify the optimal configuration of a physical system governed by a black box function $f$, namely, $x^{\star}=\arg\max_{x\in\mathcal{X}}f(x).$ The set $\mathcal{X}$ summarizes all possible system configurations, the so called search space. We assume that we can sequentially evaluate the unknown function at specific points $x$ in the search space and obtain noisy observations, $y=f(x)+\epsilon(x)$, where $\epsilon$ has a known Gaussian likelihood, which is possibly heteroscedastic. We assume that $f$ can be modeled probabilistically using a Gaussian process prior that we introduce later. Importantly, the order of the evaluations is dictated by known, potentially stochastic, dynamics modeled by a Markov chain that limits our choices of $x\in\mathcal{X}$.
@@
Figure 1: Representative task of finding pollution in a river while following the current. (a) Problem formulation: The star represents the maximizer and the arrows the Markov dynamics. (b) Objective formulation: Orange balls represent potential maximizers, with size corresponding to model uncertainty. (c) Optimization: Deploy a potentially stochastic policy that minimizes our objective.
@@ 2 Background
Because our contributions address experimental design of real-life systems by intersecting design of experiments, BayesOpt and RL, we review each of these of components. Refer to Figure 5 in Appendix A for a visual overview of how we selected the individual components for tractability of the entire problem.
@@ 2.1 Maximum Identification: Experiment Design Goal
Suppose we are in step $t$ out of $T$, then let $\mathbf{X}_{t}$ be the set of previous queries, we seek to identify new $\mathbf{X}_{\text{new}}$ that when evaluated minimize the probability of returning a sub-optimal arm at the end. For a given function draw $f$, the probability of returning a wrong maximizer $z\neq x_{f}^{\star}$ is $P(\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\geq 0|f)$. We can then consider the worst-case probability across potential maximizers, and taking expectation over $f$ we obtain a utility through an asymptotic upper-bound on the log-probability, indeed for large $T$ we obtain:
$\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\log P(\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\geq 0|f)\right]\stackrel{{\scriptstyle\cdot}}{{\leq}}-\frac{1}{2}\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\frac{(f(z)-f(x_{f}^{\star}))^{2}}{k_{\mathbf{X}_{t}\cup\mathbf{X}_{new}}(z,x_{f}^{\star})}\right]$ (3)
The expectation is on the current prior (posterior up to $\mathbf{X}_{t}$), the kernel $k$ is the posterior kernel given observations $\mathbf{X}_{t}\cup\mathbf{X}_{\text{new}}$. Since we consider the probability of an error, it is more appropriate to talk about minimizing instead of ‘maximizing the utility’ but the treatment is analogous. Further, note the intuitive interpretation of the bound: the probability of an error will be minimized if the uncertainty is small or if the values of $f(z)$ and $f(x^{\star}_{f})$ are far apart. The non-trivial distribution of $f(x^{\star})$ [29] renders the utility intractable; therefore we employ a simple and tractable upper bound on the objective (3) which can be optimized by minimizing the uncertainty among all pairs in $\mathcal{Z}$:
@@
Utility with kernel embeddings. For illustrative purposes, consider a special case where the kernel $k$ has a low rank due to existence of embeddings $\Phi(x)\in\mathbb{R}^{m}$, i.e., $k(x,y)=\Phi(x)^{\top}\Phi(y)$. Such embeddings can be, e.g., Nyström features [31] or Fourier features [32, 33]. While not necessary, these formulations make the objectives considered in this work more tractable and easier to expose to the reader. With the finite rank assumption, the random function $f$ becomes,
@@ 2.2 Markov Decision Processes
To model the transition constraints, we use the versatile model of Markov Decision processes (MDPs). We assume an environment with state space $\mathcal{X}$ and action space $\mathcal{A}$, where we interact with an unknown function $f:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$ by rolling out a policy for $H$ time-steps (horizon) and obtain a trajectory, $\tau=(x_{0},a_{0},x_{1},a_{1},...,x_{H-1},a_{H-1})$. From the trajectory, we obtain a sequence of noisy observations $y(\tau):=\{y(x_{0},a_{0}),...,y(x_{H-1},a_{H-1})\}$ s.t. $y(x_{h})=f(x_{h},a_{h})+\epsilon(x_{h},a_{h})$, where $\epsilon(x_{h},a_{h})$ is zero-mean Gaussian with known variance which is potentially state and action dependent. The trajectory is generated using a known transition operator $P(x_{h+1}|x_{h},a_{h})$. A Markov policy $\pi(a_{h}|x_{h})$ is a mapping that dictates the probability of action $a_{h}$ in state $x_{h}$. Hence, the state-to-state transitions are $P(x_{h+1},x_{h})=\sum_{a\in\mathcal{A}}\pi_{h}(a|x_{h})P(x_{h+1}|x_{h},a)$. In fact, an equivalent description of any Markov policy $\pi$ is the corresponding distribution giving us the probability of visiting a state-action pair under the policy, which we denote $d_{\pi}\in\mathcal{D}$, where
@@ 2.3 Experiment Design in Markov Chains
Notice that the utility $U$ in Eq. 6 depends on the states visited and hence states of the trajectory. In our notation, $\mathbf{X}_{t}$ will now form a set of executed trajectories. With deterministic dynamics, we could optimize over trajectories, but this would lead to an exponential blowup (i.e. $|X|^{H}$). In fact, for stochastic transitions, we cannot pick the trajectories directly, so instead we work in the space of distributions. For a given policy, through sampling, we are able to create an empirical distribution of all the state-action pairs visited during policy executions, $\hat{d}_{\pi}(x,a)$, which assigns equal mass to each state-action visited during our trajectories. This allows us to focus on the expected utility over the randomness of the policy and the environment, namely,
$\mathcal{U}(d_{\pi}):=U(\mathbb{E}_{\tau_{1}\sim\pi_{1},\dots\tau_{t}\sim\pi_{t}}[\hat{d}_{\pi}]).$ (7)
This formulation stems from Mutný et al. [17] who try to tractably solve such objectives that arise in experiment design by performing planning in MDPs. They focus on learning linear operators of an unknown function, unlike identifying a maximum, as we do here. The key observation they make is that any policy $\pi$ induces a distribution over the state-action visitations, $d_{\pi}$. Therefore we can reformulate the problem of finding the optimal policy, into finding the optimal distribution over state-action visitations as: $\min_{d_{\pi}\in\mathcal{D}}\mathcal{U}(d_{\pi})$, and then construct policy $\pi$ via marginalization. We refer to this optimization as the planning problem. The constraint $\mathcal{D}$ encodes the dynamics of the MDP.
@@ 2.4 Additional Related Works
The most relevant prior work to ours is exploration in reinforcement learning through the use of Markov decision processes as in Mutný et al. [17] and convex reinforcement learning of Hazan et al. [34], Zahavy et al. [35] which we will use to optimize the objective. Other related works are:
Pure exploration bandits objectives. Similar objectives to ours have been explored for BayesOpt. Li and Scarlett [36] use the $\mathcal{G}$-allocation variant of our objective for batch BayesOpt, achieving good theoretical bounds. Zhang et al. [37] and recently Han et al. [38] take advantage of possible maximizer sets to train localized models, while Salgia et al. [39] show that considering adaptive maximization sets yields good regret bounds under random sampling. Contrary to them, motivation and derivation in terms of a Bayesian decision rule do not appear elsewhere according to our best knowledge. We also recognize that we can relax the objective and optimize it in the space of policies.
Optimizing over sequences. Previous work has focused on planning experimental sequences for minimizing switching costs [11, 21, 40, 41] however they are only able adhere to strict constraints under truncation heuristics [20, 42, 22]. Recently, Qing et al. [43] also tackle Bayesian optimization within dynamical systems, with the focus of optimizing initial conditions. Concurrent work of Che et al. [44] tackles a constrained variant of a similar problem using model predictive control with a different goal.
@@
Non-myopic Bayesian Optimization. Look-ahead BayesOpt [48, 49, 50, 51, 52, 53, 54] seeks to improve the greedy aspect of BayesOpt. Such works also use an MDP problem formulation, however, they define the state space to include all past observations (e.g. [55, 56]). This comes at the cost of simulating expensive integrals, and the complexity grows exponentially with the number of look-ahead steps (usually less than three steps). Our work follows a different path, we maintain the greedy approach to control computational efficiency (i.e. by optimizing over the space of Markovian policies), and maintain provable and state-of-art performance. Even though the optimal policy through non-myopic analysis is non-Markovian, in Sec. 4, we show that adaptive resampling iteratively approximates this non-myoptic optimal policies in a numerically tractable way via receeding horizon planning. In our experiments we comfortably plan for over a hundred steps.
@@ 3 Transition Constrained BayesOpt
Motivated by our practical experiments with chemical reactors, we distinguish two different types of feedback. With episodic feedback we can be split the optimization into episodes. At the end of each episode of length $H$, we obtain the whole set of noisy observations. On the other hand, instant feedback is the setting where we obtain a noisy observation immediately after querying the function. Asynchronous feedback describes a mix of the previous two, where we obtain observations with unspecific a delay.
@@ 3.1 Expected Utility for Maximizer Identification
where $\mathbf{V}(d_{\pi})=\left(\sum_{x,a\in\mathcal{X}\times\mathcal{A}}\frac{d_{\pi}(x,a)\Phi(x,a)\Phi(x,a)^{\top}}{\sigma^{2}(x,a)}+\frac{1}{TH}\mathbf{I}\right)$. The variable $d_{\pi}(x,a)$ is a state-action visitation, $\Phi(x)$ are e.g. Nyström features of the GP. We prove that the function is additive in terms of state-action pairs in Lemma D.1 in Appendix D, a condition required for the expression as a function of state-action visitations [17]. Additionally, by rewriting the objective in this form, the dependence and convexity with respect to the state-action density $d_{\pi}$ becomes clear as it is only composition of a linear function with an inverse operator. Also, notice that the constraint set is a convex polytope. Therefore we are able to use convex optimization to solve the planning problem (see Sec. 4).
@@
In the Eq. (8) we need to approximate a maximum over all input pairs in $\mathcal{Z}$. While this can be enumerated in the discrete case without issues, it poses a non-trivial constrained optimization problem when $\mathcal{X}$ is continuous. As an alternative, we propose approximating the set $\mathcal{Z}$ using a finite approximation of size $K$ which can be built using Thompson Sampling [57, 58] or through maximization of different UCBs for higher exploitation (see Appendix E.1). In Appendix E.5, we numerically benchmark reasonable choices of $K$, and show that the performance is not significantly affected by them.
@@
  Input: Procedure for estimating sets of maximizers, initial point $x_{0}$, initial set of maximizer candidates $\mathcal{Z}_{0}$
@@
           $y_{t,h+1}=f(x_{t,h+1})+\epsilon_{t,h}$ // obtain observation
           $\mathcal{GP}_{t,h},\ \mathcal{Z}_{t,h}\leftarrow\text{Update}(\mathbf{X}_{t,h},\mathbf{Y}_{t,h})$ // update model and maximizer candidate set
@@
        $\mathbf{Y}_{t,H}=f(\mathbf{X}_{t,H})+\vec{\epsilon}_{t,:}$ // obtain observations
        $\mathcal{GP}_{t+1,:},\ \mathcal{Z}_{t+1,:}\leftarrow\text{Update}(\mathbf{X}_{t,H},\mathbf{Y}_{t,H})$ // update model and maximizer candidate set